OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -9 - 5t. Find the instantaneous velocity at t = 4 by finding the derivative

ganeshie8 (ganeshie8):

yes so lets find the derivative

ganeshie8 (ganeshie8):

\(\large s(t) = -9-5t\) \(\large v(t) = s'(t) = ?\)

OpenStudy (anonymous):

ummm how do I work this?

ganeshie8 (ganeshie8):

\[\large \dfrac{d}{dx}(x^n) = nx^{n-1}\] seen this formula before ?

OpenStudy (anonymous):

I think so

ganeshie8 (ganeshie8):

and you may be knowing that derivative of a constant is 0 : \[\large \dfrac{d}{dx}(c) = 0\]

OpenStudy (anonymous):

Okay

ganeshie8 (ganeshie8):

\[\large v(t) = \dfrac{d}{dt}(s(t)) = \dfrac{d}{dt}(-9-5t)\]

ganeshie8 (ganeshie8):

whats the derivative of "-9" ?

OpenStudy (anonymous):

0?

ganeshie8 (ganeshie8):

yes! what about the derivative of "-5t" ?

ganeshie8 (ganeshie8):

"-5t" looks like "y=mx" right ? so derivative is just "-5"

ganeshie8 (ganeshie8):

\[\large v(t) = \dfrac{d}{dt}(s(t)) = \dfrac{d}{dt}(-9-5t) \] \[\large = 0-5 \] \[\large = -5 \]

ganeshie8 (ganeshie8):

\[\large v(t) = -5 \]

ganeshie8 (ganeshie8):

see if that makes more or less sense..